OpenStudy (anonymous):

trying to find the second derivative in this question......3x-3 over 2x+4. the Vertical Asy is -2 and the Horizontal Asy is 3/2...the xint is (1,0) and the yint is (0,-3/4) please give me step buy step

OpenStudy (anonymous):

can you find the first derivative? because that is the only hard part

OpenStudy (anonymous):

quotient rule give \[\frac{d}{dx}\frac{3x-3}{2x+4}=\frac{(2x+4)\times 3-(3x+3)\times 2}{(2x+4)^2}\]

OpenStudy (anonymous):

yes i have the first one all ready

OpenStudy (anonymous):

so you end up with \[\frac{9}{(2x+4)^2}=9(2x+4)^{-2}\] now second should be easy

OpenStudy (anonymous):

\[-18(2x+4)^{-3}\times 2=-36(2x+4)^{-3}=-\frac{36}{(2x+4)^3}\]

OpenStudy (anonymous):

you can cancel a little because \[(2x+4)^3=(2(x+2))^3=8(x+2)^3\]

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

ok which step? did you get the first derivative?

OpenStudy (anonymous):

i got (2x+4)-2(3x-3) over (2x+4)^2

OpenStudy (anonymous):

then you have to do the algebra in the numerator

OpenStudy (anonymous):

should be \[(2x+4)3-2(3x-3) \]

OpenStudy (anonymous):

in front of the 2x+4 there is suppose to be a 3

OpenStudy (anonymous):

yes it is the denominator times the derivative of the numerator as the first term

OpenStudy (anonymous):

minus the numerator times the derivative of the denominator

OpenStudy (anonymous):

if you multiply out the x terms add up to zero and you get 9

OpenStudy (anonymous):

so 18 over (2x+4)^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(2x+4)3-2(3x-3)\] \[6x+12-6x+6=18\] yes

OpenStudy (anonymous):

my equation is 3(2x+4)-2(3x-3)

OpenStudy (anonymous):

yes multiply out and you will get 18

OpenStudy (anonymous):

ok then what do i do to get the second answer

OpenStudy (anonymous):

you have \[\frac{18}{(2x+4)^2}\] so rather than using the quotient rule again, rewrite in exponential form as \[18\times (2x+4)^{-2}\] and use the power rule (and the chain rule)

OpenStudy (anonymous):

ok so multiple it out?

OpenStudy (anonymous):

power rule right? \[\frac{d}{dx}x^n=nx^{n-1}\]

OpenStudy (anonymous):

i made a mistake, sorry

OpenStudy (anonymous):

\[\frac{d}{dx}18(2x+4)^{-2}=-2\times 18(2x+4)^{-2-1}\times 2\] \[=-72(2x+4)^{-3}=-\frac{72}{(2x+4)^3}\] but you can simplify this

OpenStudy (anonymous):

because \[(2x+4)^3=(2(x+2))^3=8(x+2)^3\] and \[\frac{72}{8}=9\]

OpenStudy (anonymous):

so your "final answer" will be \[-\frac{9}{(x+2)^3}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw